OpenStudy (anonymous):

PLEASE HELP!!! Write the equation of the hyperbola where the foci are at (-5,4) and (-5,-10), and a=5.

OpenStudy (neer2890):

centre at (0,0)??

OpenStudy (imstuck):

No the center is at (-5, -3)

OpenStudy (anonymous):

Center is mid point of foci i.e (-5,-3)

OpenStudy (anonymous):

how did you find the midpoint of foci?

OpenStudy (anonymous):

@ShailKumar

OpenStudy (anonymous):

Mid point of two points (x1,y1) and (x2,y2) is ( (x1+x2)/2 , (y1+y2)/2 )

OpenStudy (anonymous):

oh i see for the hyperbola is the center the same as h,k then?

OpenStudy (imstuck):

yes

OpenStudy (anonymous):

so then i have (x+5)^2+(y+3)^2=1 but its over a^2 and b^2 if we have a how do we find b? how can we find c to do the a^2+b^2=c^ thing?

OpenStudy (imstuck):

This is a hyperbola of the form\[\frac{ (y-k)^{2} }{ a ^{2} }-\frac{ (x-h)^{2} }{ b ^{2} }=1\]

OpenStudy (anonymous):

yes. But in this problem you cannot apply standard for of hyperbola. Standard forms are meant for either horizontal or vertical hyperbola. This is not the case here.

OpenStudy (anonymous):

*form

OpenStudy (anonymous):



OpenStudy (imstuck):

If this isn't a horizontal or a vertical hyperbola, then what kind is it?

OpenStudy (anonymous):

Oh yes, its a vertical hyperbola

OpenStudy (anonymous):

how do you know its a vertical hyperbola and why is that significant?

OpenStudy (imstuck):

C is equal to the number of units up from the vertex the foci is. Here the foci is 7 units up and 7 units down from the vertex. So c = 7 c^2 = 49. a^2 + b^2= c^2...now fill in what you have. 49 = 25 + b^2 or 49 - 25 = b^2...24 = b^2. There's your b^2 for your formula.

OpenStudy (anonymous):

so b is square root 24?

OpenStudy (anonymous):

wai b^2 is just 24 and b is square root 24 right?

OpenStudy (imstuck):

This is vertical because the foci lie on an "x = " line. The foci lie on the x = -5 line so it is a vertical hyperbola. It's important because the x^2 and the y^2 terms move according to which type of hyperbola it is. Like in an ellipse, the a^2 and b^2 move and the a^2 is always under the major axis. In hyperbolas, the y^2 coming first indicates the foci are on the y axis. The x^2 coming first indicates the foci are on the x axis. And yes, b is the square root of 24.

OpenStudy (imstuck):

\[\frac{ (y+3)^{2} }{ 25 }-\frac{ (x+5)^{2} }{ 24 }=1\]

OpenStudy (anonymous):

does it matter if you do the (y+3) or (x+5) first?

OpenStudy (imstuck):

Remember the y axis is a vertical line. Look at a graph. And an x = something line is also a vertical line. But it's easier to remember the y^2 because the foci lie along the y axis, up and down, not side to side. Yes it matters very much if you do the (y + 3)^2 first; this indicates the way the hyperbola lie on the graph...it DEFINES the type of hyperbola it is.

OpenStudy (imstuck):